if A is a positive acute angle and tan A = cot 30 then what is the value of tan A
Answer by jim_thompson5910(35256)

tan(A) = cot(30)

tan(A) = cos(30)/sin(30)

tan(A) = (sqrt(3)/2)/(1/2)

tan(A) = (sqrt(3)/2)*(2/1)

tan(A) = sqrt(3)


Answer:
tan(A) = sqrt(3)